FIND THE WITDTH OF A RECTANGLE WHOSE WIDTH IS 1 MORE THAN TWICE ITS LENGTH AND WHOSE PERIMERTER IS 26

Answer by algebrahouse.com(1659)

FIND THE WITDTH OF A RECTANGLE WHOSE WIDTH IS 1 MORE THAN TWICE ITS LENGTH AND WHOSE PERIMERTER IS 26.

x = length
2x + 1 = width {width is one more than twice the length}

2(x) + 2(2x + 1) = 26 {two times width plus two times length equals perimeter}
2x + 4x + 2 = 26 {used distributive property}
6x + 2 = 26 {combined like terms}
6x = 24 {subtracted 2 from both sides}
x = 4 (length) {divided both sides by 4}
2x + 1 = 9 (width)
www.algebrahouse.com